A river located in Antalya is open to all fishermen so they can fishing freely. They shoud pay 10 dollar for a boat for fishing. The equation is f(n)= 90n^1/3. n is number of boats. Equation gives total number of sea bass caught and each boat catches f(b)/ b kg sea bass. The price of a sea bass is 1 dollar per kg.

Can you find the efficient number of boats

Can you find per boat tax would statisy efficiency .
Answer #1

Total cost for n boats = 10n

Total Revenue = 90(n)^1/3

Total profit = 90(n)^1/3 - 10n

for maximum : dP/dn =0

=> 90*(1/3)/n^2/3 -10 = 0

=> n^2/3 = 3

=> n = 3^1.5 = 3*1.7 = 5(approx)

efficient value = 90(5.1)^1/3 - 10(5) = 90*1.72 - 50 = 105

2) let tax = t

Total profit per boat = 90/n^2/3 -10

after tax this amount should be 90/n^2/3 - 10 - t

this amount should be 105/5 = 21

90/n^2/3 -t = 31

t = 90/(n^2/3) -21